OpenStudy (anonymous):

can someone check this i got 3, -1 solve each system by elimination Please help me...Reward given :) -x+2/5=-3/5y 3y=-18/11x+51/11

hero (hero):

Oh wow :)

hero (hero):

At least you attempted this one it seems

OpenStudy (zarkon):

if the equations are \[-x+\frac{2}{5}=\frac{-3}{5}y\] and \[3y=-\frac{18}{11}x+\frac{51}{11}\] then no

OpenStudy (anonymous):

then no what

OpenStudy (zarkon):

3, -1 is not the solution

OpenStudy (anonymous):

really ?

OpenStudy (zarkon):

x=1,y=1